### v1

# A proof of Alexander’s conjecture on an inequality of Cassels

Abstract.
Let $z_{1},\dots,z_{n}$ be complex numbers with $|z_{j}|\leq\rho$, where $\rho>1$. Cassels proved that, under an additional restriction on $\rho$, the inequality
$\prod_{j\neq k}\bigl|1-\overline{z_{j}}z_{k}\bigr|\leq\left(\frac{\rho^{2n}-1}{\rho^{2}-1}\right)^{\!n}$
holds. In a subsequent note, Alexander conjectured that this inequality is in fact valid without any restriction on $\rho$. In this paper, we confirm Alexander’s conjecture.
Key words and phrases:
Cassels inequality, Alexander conjecture, Schur inequality, complex inequalities, extremal configurations
2020 Mathematics Subject Classification:
30C80, 26D15
1. Introduction
Let $z_{1},\dots,z_{n}$ be complex numbers in the unit disc. A classical inequality, going back to Schur [1], states that
(1) $\prod_{j\neq k}\bigl|1-\overline{z_{j}}z_{k}\bigr|\leq n^{n},$
with equality if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
In connection with a problem of Schinzel and Zassenhaus, Cassels [2] extended (1) with the following:
Theorem 1 (Cassels).
Let $\rho>1$ and let $z_{1},\dots,z_{n}$ be complex numbers in the disc $\{z:|z|\leq\rho\}$. Suppose
(2) $\cos(\pi/n)\leq\frac{\rho^{2}}{\rho^{4}-\rho^{2}+1}.$
Then
(3) $\prod_{j\neq k}\bigl|1-\overline{z_{j}}z_{k}\bigr|\leq\left(\frac{\rho^{2n}-1}{\rho^{2}-1}\right)^{\!n},$
with equality if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
Alexander [3] observed that condition (2) may be replaced by the weaker assumption
$\cos(\pi/n)\leq\frac{2\rho^{2}}{\rho^{4}+1},$
and conjectured that (3) should in fact hold for all $\rho>1$, without assuming (2).
In a recent work, Dubickas [4] revisited Cassels’ inequality by observing that, for $z_{j}=\rho\,\omega_{j}$ with $|\omega_{j}|=1$, one has
$\prod_{j\neq k}\bigl|1-\overline{z_{j}}z_{k}\bigr|=\rho^{n(n-1)}\prod_{j<k}\left((\rho-\rho^{-1})^{2}+|\omega_{j}-\omega_{k}|^{2}\right).$
He conjectured that the elementary symmetric functions of the quantities $|\omega_{j}-\omega_{k}|^{2}$ attain their maximum for the vertices of a regular $n$-gon inscribed in the unit circle, and proved this for the elementary symmetric functions of degree 11 through 44. His conjecture implies Alexander’s.
In the present paper, we confirm Alexander’s conjecture. Our approach relies on an auxiliary additive inequality and a monotonicity argument.
2. Main result
Theorem 2.
Let $\rho>1$ and let $z_{1},\dots,z_{n}$ be complex numbers in the disc $\{z:|z|\leq\rho\}$. Then the following inequality holds:
(4) $\prod_{j\neq k}|1-\overline{z_{j}}z_{k}|\leq(\rho^{2n}-1)^{n}(\rho^{2}-1)^{-n}.$
Equality is attained if and only if the $z_{j}$ are the vertices of a regular $n$-gon inscribed in the circle of radius $\rho$.
3. An auxiliary inequality
Theorem 3.
Let $z_{1},\dots,z_{n}$ on the unit disc and denote $\Lambda=(-1)^{n}\prod_{j=1}^{n}z_{j}$. Then we have the following inequality:
(5) $\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}}\geq\frac{n^{2}}{1-|\Lambda|^{2}}.$
Equality is attained if and only if the $z_{j}$ are the $n$ roots of the polynomial $\zeta^{n}+\Lambda$.
Corollary 4.
Let $\omega_{1},\dots,\omega_{n}$ on the unit circle and $a\in(0,1)$. Then the following inequality holds:
(6) $\sum_{j,k}\frac{1}{1-a\overline{\omega_{j}}\omega_{k}}\geq\frac{n^{2}}{1-a^{n}}.$
Equality is attained if and only if the $\omega_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
4. Proof of the main theorem
Proof of Theorem 2.
Fix an index $k$. The map
$z_{k}\longmapsto\prod_{j\neq k}(1-\overline{z_{j}}z_{k})$
is holomorphic in the disc $\{|z_{k}|\leq\rho\}$. By the maximum modulus principle, the quantity $\prod_{j\neq k}|1-\overline{z_{j}}z_{k}|$ is maximized when $|z_{k}|=\rho$. Iterating this argument for $k=1,\dots,n$ (applied to the factors involving $z_{k}$ in the product) shows that
(7) $\prod_{j\neq k}|1-\overline{z_{j}}z_{k}|\leq\max_{|\omega_{1}|=\cdots=|\omega_{n}|=1}\prod_{j\neq k}|1-\rho^{2}\overline{\omega_{j}}\omega_{k}|,$
and equality can occur only if $|z_{1}|=\cdots=|z_{n}|=\rho$.
Setting $z_{j}=\rho\,\omega_{j}$ with $|\omega_{j}|=1$, (4) is equivalent to
(8) $\prod_{j,k}\bigl|\rho^{-2}-\overline{\omega_{j}}\omega_{k}\bigr|\leq(1-\rho^{-2n})^{n}.$
We fix $\omega_{1},\dots,\omega_{n}$ in the unit circle and we define, for $a\in[0,1)$,
$g_{\omega}(a)=n\ln(1-a^{n})-\sum_{j,k}\ln|a-\overline{\omega_{j}}\omega_{k}|.$
Using $\frac{d}{da}\log|a-\zeta|=\Re\frac{1}{a-\zeta}$ and the fact that $\sum_{j,k}\!\left(\frac{1}{a-\overline{\omega_{j}}\omega_{k}}\right)$ is real, we compute
[…]
It follows from Corollary 4 that $g_{\omega}^{\prime}(a)\geq 0$ for all $a\in[0,1)$. Since $g_{\omega}(0)=0$, we conclude that $g_{\omega}(\rho^{-2})\geq 0$, which proves (8) and hence the desired inequality (4).
Now assuming that $z_{1},\dots,z_{n}$ realise equality in (4), they also realise equality in (7). Hence $z_{j}=\rho\omega_{j}$ with $|\omega_{j}|=1$ for $j=1,\dots,n$ and $g_{\omega}(\rho^{-2})=0$. Since $g_{\omega}$ is nondecreasing and $g_{\omega}(0)=g_{\omega}(\rho^{-2})=0$, it follows that $g_{\omega}$ is constant on $0,\rho^{-2}]$, hence $g_{\omega}^{\prime}(\rho^{-2})=0$, that is,
$\sum_{j,k}\frac{1}{1-\rho^{-2}\overline{\omega_{j}}\omega_{k}}=\frac{n^{2}}{1-\rho^{-2n}}.$
By Corollary 4, the $\omega_{j}$ are the vertices of a regular $n$-gon inscribed in the unit circle.
Conversely, if the $z_{j}$ are the vertices of a regular $n$-gon, then
$\prod_{j=1}^{n}(\zeta-z_{j})=\zeta^{n}+\Lambda,\qquad|\Lambda|=\rho^{n}.$
Hence
$\prod_{j,k}\bigl|1-\overline{z_{j}}z_{k}\bigr|=\rho^{n}\prod_{j=1}^{n}\bigl|\overline{z_{j}}^{-n}+\Lambda\bigr|=\rho^{n}\prod_{k=1}^{n}\bigl|-\overline{\Lambda}^{-1}+\Lambda\bigr|=(\rho^{2n}-1)^{n}.$
5. Proof of Theorem 3
Proof.
Fix $z_{1},\dots,z_{n}$ on the unit disc and define
$f(\zeta):=\sum_{j=1}^{n}\frac{1}{1-\overline{z_{j}}\zeta},$
and the finite Blaschke product
$B(\zeta):=\prod_{j=1}^{n}\frac{\zeta-z_{j}}{1-\overline{z_{j}}\zeta}.$
Both are holomorphic in a neighborhood of the unit disc.
Expand
$|f(e^{it})|^{2}=\sum_{j,k}\frac{1}{(1-\overline{z_{j}}e^{it})(1-z_{k}e^{-it})}.$
Then
$\frac{1}{2\pi}\int_{0}^{2\pi}\frac{dt}{(1-\overline{z_{j}}e^{it})(1-z_{k}e^{-it})}=\frac{1}{2\pi i}\oint_{|\zeta|=1}\frac{d\zeta}{(1-\overline{z_{j}}\zeta)(\zeta-z_{k})}=\frac{1}{1-\overline{z_{j}}z_{k}},$
by the Cauchy integral formula. Summing over $j,k$ gives
(9) $\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}}=\frac{1}{2\pi}\int_{0}^{2\pi}|f(e^{it})|^{2}dt.$
For all $t\in[0,2\pi]$, since $|B(e^{it})|=1$, we have
$\left|1-\Lambda\overline{B}(e^{it})\right|^{2}=1+|\Lambda|^{2}-\overline{\Lambda}B(e^{it})-\Lambda\overline{B}(e^{it}).$
Integrating and using Cauchy integral formula $\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,dt=B(0)=\Lambda$ yields
(10) $\begin{split}\frac{1}{2\pi}\int_{0}^{2\pi}\left|1-\Lambda\overline{B}(e^{it})\right|^{2}dt=1-|\Lambda|^{2}.\end{split}$
Using twice again Cauchy integral formula,
(11) $n=f(0)=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\,dt,$
and
(12) $\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,\overline{f}(e^{it})\,dt=\sum_{j=1}^{n}\frac{1}{2\pi}\int_{0}^{2\pi}B(e^{it})\,\frac{e^{it}}{e^{it}-z_{j}}\,dt=\sum_{j=1}^{n}B(z_{j})=0.$
Now, we multiply the complex conjugate of (12) by $\Lambda$ and we subtract it from (11) to find
(13) $\begin{split}n&=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\,dt=\frac{1}{2\pi}\int_{0}^{2\pi}f(e^{it})\left(1-\Lambda\overline{B}(e^{it})\right)\,dt.\end{split}$
Applying Cauchy–Schwarz inequality to the integral in (13) and using (9) and (10) gives the desired inequality
(14) $\begin{split}n^{2}&\leq\Big(\frac{1}{2\pi}\int_{0}^{2\pi}|1-\Lambda\overline{B(e^{it})}|^{2}\,dt\Big)\Big(\frac{1}{2\pi}\int_{0}^{2\pi}|f(e^{it})|^{2}\,dt\Big)\\ &=(1-|\Lambda|^{2})\sum_{j,k}\frac{1}{1-\overline{z_{j}}z_{k}}.\end{split}$
Moreover, equality in (14) holds if and only if
(15) $f=c(1-\overline{\Lambda}B),$
for some constant $c$. Evaluating in 0, we find that
$c=\frac{n}{1-|\Lambda|^{2}}.$
Let $e_{m}=e_{m}(z_{1},\dots,z_{n})$ be the elementary symmetric polynomials:
$e_{m}=\sum_{1\leq j_{1}<\cdots<j_{m}\leq n}z_{j_{1}}\cdots z_{j_{m}},$
Put
$\begin{split}P(\zeta)&=\prod_{j=1}^{n}(\zeta-z_{j})=\sum_{m=0}^{n}(-1)^{\,n-m}e_{n-m}\,\zeta^{m}\\ Q(\zeta)&=\prod_{j=1}^{n}(1-\overline{z_{j}}\zeta)=\sum_{m=0}^{n}(-1)^{m}\overline{e_{m}}\,\zeta^{m},\\ R(\zeta)&=\sum_{j=1}^{n}\prod_{k\neq j}(1-\overline{z_{k}}\zeta)=\sum_{m=0}^{n-1}(-1)^{m}(n-m)\,\overline{e_{m}}\,\zeta^{m},\end{split}$
where the last identity follows by counting: each monomial of degree $m$ in the $\overline{z_{j}}$ is omitted by exactly $m$ indices and therefore appears exactly nmn-m times.
We have
$B=\frac{P}{Q},\ f=\frac{R}{Q}.$
Multiplying (15) by $(1-|\Lambda|^{2})Q$, we obtain
[…]
Identifying the coefficients in front of $\zeta^{m}$ for $m=1,\dots,n-1$, we get
(16) $(n-(1-|\Lambda|^{2})(n-m))(-1)^{m}\overline{e_{m}}=n\overline{\Lambda}(-1)^{n-m}e_{n-m}.$
Also replacing $m$ by nmn-m and taking the conjugate,
(17) $(n-(1-|\Lambda|^{2})m)(-1)^{n-m}e_{n-m}=n\Lambda(-1)^{m}\overline{e_{m}}.$
Identities (16) and (17) yield to
$\begin{split}(n-(1-|\Lambda|^{2})m)(n-(1-|\Lambda|^{2})(n-m))\overline{e_{m}}&=n^{2}|\Lambda|^{2}\overline{e_{m}}\\ \Leftrightarrow(1-|\Lambda|^{2})^{2}m(n-m)\overline{e_{m}}&=0.\end{split}$
We deduce that $e_{m}=0$ for $m=1,\dots,n-1$ and consequently that
$P(\zeta)=\zeta^{n}+(-1)^{n}e_{n}=\zeta^{n}+\Lambda.$
Therefore $z_{1},\dots,z_{n}$ are precisely the roots of $\zeta^{n}+\Lambda$.
Conversely, if $z_{1},\dots,z_{n}$ are the roots of $\zeta^{n}+\Lambda$, then $e_{m}=0$ for $1\leq m\leq n-1$ and the above coefficient comparison can be reversed to obtain (15), hence equality in (14). ∎